OpenStudy (star):

let h(x) = (sin(x^2) - x^2 cos (x))/(cos(x^2)-e^(sinx)+ln(1-x)+ax^3) what is the limit as x -> 0 as a function of a I will type the equation in a reply below!

OpenStudy (star):

\[\frac{ \sin (x^2) - x^2 \cos(x) }{ \cos(x^2) - e^{sinx} +\ln(1-x) + ax^3 }\]